N a triangle ABC if B= 75 and C = 45 then the ratio of the sides is

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Hence angle A = 180 -(75+45) = 60°
 
Now using sine rule
a/ sinA = b/ sin B = c/ sinC = k
 
Hence
a = k sinA = k* (sqrt(3)/2)
b = k sin B = k* (sqrt(3) +1)/2 sqrt(2)
c = k sinC = k* (1/sqrt(2))
 
Now I hope you can find the ratio.
If you can't, please feel free to ask again.
 
Regards
Arun (askIITians forum expert)